OpenStudy (anonymous):

What are the steps to factor 36x - 49x^3?

OpenStudy (john_es):

First extract common factor x, \[x(36-49x^2)\] Then solve the second product, \[36-49x^2=0\Rightarrow x=\pm\frac{6}{7}\] So, we have solve \[x(x-6/7)(x+6/7)=0\] Manipulating, to express in factors \[x(7x-6)(7x+6)\]